### v1

# ROI-GS: Interest-based Local Quality 3D Gaussian Splatting

Abstract
We tackle the challenge of efficiently reconstructing 3D scenes with high detail on objects of interest. Existing 3D Gaussian Splatting (3DGS) methods allocate resources uniformly across the scene, limiting fine detail to Regions Of Interest (ROIs) and leading to inflated model size. We propose ROI-GS, an object-aware framework that enhances local details through object-guided camera selection, targeted Object training, and seamless integration of high-fidelity object of interest reconstructions into the global scene. Our method prioritizes higher resolution details on chosen objects while maintaining real-time performance. Experiments show that ROI-GS significantly improves local quality (up to PSNR), while reducing overall model size by of baseline and achieving faster training for a scene with a single object of interest, outperforming existing methods.
I Introduction
Reconstructing complex 3D scenes with high-quality details objects of interest remains a long-standing goal in computer vision and graphics. Like human vision, which naturally focuses on salient objects or depth-varying areas while leaving peripheral regions less resolved, this selective fidelity enables efficient perception. Similarly, computational methods benefit from concentrating resources on semantically or geometrically important regions. This capability is particularly vital in industrial applications and cultural heritage contexts. Here, our work aims to create 3D digital twins of cultural heritage sites, facilitating preservation [1], virtual exploration [2], and interactive museum experiences [3], thus supporting archiving, education, and public participation while minimizing physical handling and degradation risks.
Recent advances in 3D Gaussian Splatting (3DGS) [4] have rapidly gained traction as a powerful alternative for neural scene representation, offering explicit scene models, real-time rendering, and state-of-the-art visual quality. However, its explicit nature comes at the cost of a large memory footprint that can become prohibitive in large-scale scenes that require high detail in the internal contents of interest, as reported in [4, 5, 6]. This challenge is exacerbated by uniform scene reconstruction in most existing 3DGS methods, which allocates equal resources to all regions regardless of their importance. Prior work on level of detail (LOD) Gaussian Splatting [7, 6, 8, 9, 10, 11, 12, 13, 14] or scene chunking [15, 16] improves overall quality and scalability, but does not account for the presence of content of interest within the scene. Efforts in object-centric [17] and segmentation-aware [18, 19, 20] modeling focus capacity on individual scene elements. While these methods enable precise object handling, they require coherent object masks across images and do not adaptively select or schedule views based on objects of interest.
We address these gaps with ROI-GS, an object-aware 3DGS framework that reconstructs user-specified objects of interest using optimized image subsets and reintegrates them into the overall scene model. Our pipeline includes a scene decomposition module based on object-specific automatic camera selection, enabling independent training of scene components. By focusing computational resources and details on user-selected ROIs, the approach achieves fine-grained ROI optimization without significantly increasing the overall memory footprint. Leveraging the explicit nature of 3DGS representation, we seamlessly replace lower-LOD Gaussians in the base scene with refined ROI Gaussians, producing a unified model that combines global context at moderate quality with high-quality local regions, all rendered in real-time.
The main contributions of this work are: (i) we introduce region-aware reconstruction into the 3DGS paradigm through object-guided view selection and targeted Object training; (ii) we improve reconstruction quality using an advanced image selection strategy based on model optimization that maximizes ROI coverage and occupancy; and (iii) we propose a simple yet effective composition strategy that integrates high-fidelity Object Gaussians into the overall scene model. Together, these contributions enable ROI-GS to enhance object-level detail in complex scenes without sacrificing real-time rendering performance under a practical number of ROIs. Empirically, our method achieves up to a improvement in PSNR within ROI, and reduces overall model size, in this work, composed from a scene trained with iterations along with a single object of interest, to compared to baselines.
II Method
Our ROI-GS pipeline builds on the standard 3D Gaussian Splatting framework, which models the scene as a set of anisotropic 3D Gaussians. It takes as input a set of 2D color images, along with camera poses, intrinsic parameters, and a sparse point cloud, producing highly realistic real-time renderings with enhanced LOD in regions of interest (cf. Figure 2).
II-A Object-Focused Camera Selection.
We propose a two-step ROI-focused camera selection strategy. First, each user-specified ROI is defined by an axis-aligned bounding box (AABB) enclosing the object of interest. In the initial filtering step, we select all views that observe the object by identifying cameras that capture at least one 3D keypoint within the bounding box, based on point-to-image visibility from SfM data. This first filtering significantly narrows down the set of candidate images. However, using all these views to train an Object-GS may degrade object detail due to low resolution from distant cameras and also reduce computational efficiency. To address this, we introduce a refinement step that assesses and selects the most informative and optimal views for high-quality object reconstruction. We initially used a simple approach based on static criteria, such as the distance between the ROI and viewpoints, the projected area of the ROI’s bounding box on 2D images, or the number of reconstructed keypoints within the box and visible to each camera. While effective in identifying views rich in ROI information, these criteria could introduce spatial bias, favoring regions with dense camera clusters and resulting in incomplete or unbalanced coverage from diverse perspectives. To mitigate this, we adopt an advanced selection strategy based on model-driven optimization. Prior works on View selection in GS [21, 22, 23, 24] reduce image redundancy and maximize scene coverage by selecting globally informative views, but they do not account for object-specific camera relevance. Our advanced selection approach relies on ActiveInitSplat [22], with adaptations for object-focused optimization. The views are ordered based on the ROI’s 3D point quality score: the next view is selected as the one that maximizes the improvement of this score. Specifically, this refinement selection module considers two terms of the sparse point distribution: the point density and voxel occupancy within the ROI bounding box to identify any sparsely covered areas or angles. A Gaussian Process-based (GP) model is used to guide this selection. Furthermore, by incorporating the static criteria mentioned above as parameters, we achieve more accurate and consistent camera selection for the target object, leading to better-aligned Gaussian optimization and ultimately higher-fidelity ROI reconstructions. The original model uses input viewpoint parameters ( for position and for orientation). Our extended version adds static parameters (distance, projected AABB area, and number of keypoints), totaling inputs. We compare the -params and -params versions in Section III-C.
II-B GS Training
Once the relevant images are selected and grouped, we train separate high-quality Gaussian Splatting models for each object (Object-GS) and one for the global scene (Scene-GS). The Scene-GS model is first trained for iterations using all scene images, excluding the subset of close-up views reserved for ROI training. To maintain consistency, a portion of the ROI images can optionally be retained in the Scene-GS training to give the global model a coarse understanding of the object’s geometry. In our experiments, of the ROI selected images are included in Scene-GS training.
In early experiments, we attempted to initialize Object-GSs from the sparse SfM point cloud. However, neglecting the global representation of the trained Scene-GS model proved inefficient. The next implementation idea was to initialize only Scene-GS Gaussians restricted to the ROI’s volume, and continue optimizing only within this region. However, the lack of out-of-the-box context led to floating artifacts, as the model hallucinated missing geometry beyond the object of interest, significantly degrading reconstruction quality. To overcome this, we initialize each Object-GS directly from the full Scene-GS model. This approach provides both contextual grounding and consistently yields higher-quality object reconstructions, making it a more effective and efficient starting point for ROI refinement. Each Object-GS is then trained independently using only the subset of selected images focused on the target object. The object training follows standard 3DGS optimization for iterations to ensure sufficient reconstruction quality. Optimization is applied to the entire image, updating Gaussians even in the vicinity of the object box and the background, to prevent floating artifacts within the ROI. However, the Gaussian densification process is confined to the ROI volume for iterations, keeping detail enhancement limited within the target region. Additionally, a pruning strategy is applied during training to remove non-contributing Gaussians, many of which were inherited from the full Scene-GS initialisation, thereby reducing the memory footprint of the Object-GS model. Focusing on a bounded region with high-resolution inputs allows for significantly finer LOD in this region than the global model. Object-GS can thus capture high-fidelity representations, including geometric details and texture appearance that could otherwise be blurred or undersampled in a uniformly trained entire-scene model.
II-C Scene–Objects Composition with Gaussians.
For the inference step, we integrate the Object-GSs models back into the global Scene-GS. This composition is simple and efficient thanks to the explicit nature of 3DGS scene representation. For each object, we replace the original Scene-GS Gaussians that lie inside the object’s bounding box with the corresponding fine details Object-GS ones. Since all Gaussians share a common coordinate system from SfM, the replacement is seamless and co-registered. The result is a combined Gaussian set representing the entire scene, where each ROI is now filled with its high-quality Gaussians.
III Experiments
In this Section, we discuss the implementation details of our ROI-GS method and present the method’s performance in enhancing object-level reconstruction quality in a complex real-world scene.
III-A Experimental setup.
We conduct experiments on the Bureau scene of Hôtel de la Marine dataset, which we captured at the cultural heritage site Hôtel de la Marine in Paris111https://www.hotel-de-la-marine.paris/. It consists of images at resolution, which walk through the entire hotel working room, including close-ups of objects on a desk, such as books, coffret, or candlesticks. To reduce storage and computational costs, the image resolution was reduced by a factor of 4, which is equivalent to pixels.
Due to the lack of space and to enable fair comparison and objective performance evaluation, in this experiments we only consider a single object of interest. However, the method is scalable by design as objects can be handled independently. The object chosen here is a coffret on the desk. The number of images that have at least a visible keypoint in the box is . From these images, are retained as a test set to evaluate the performance of the model. Our implementation is based on the 3DGS framework [4]. All trainings are performed using a single H100 GPU on a DGX workstation.
The baseline models used for comparison are: (1) Full Scene, which is trained on all images in the scene, and (2) Full Object, which is trained on all images where the object of interest is visible. Both models are optimized for 50K iterations to ensure maximizing the high-quality scene reconstruction.
We evaluated the Object-Focused Camera Selection technique on two versions of the Advanced GP-based model. In our experiments, the typical number of selected cameras for ROI training is out of all ROI-visible images. Selecting more images was found to slightly degrade reconstruction quality, as distant or less informative views may lead to reduced resolution due to insufficient object resolution. Both the original GP-based model (Advanced 6-params GP) and its extended version (Advanced 9-params GP), described in Section II-A, select the first cameras while preserving their order. Unlike the default 3DGS pipeline, where camera order is irrelevant, the advanced selection strategy determines an order for selected views. Therefore, during training, we retain this order instead of applying random shuffling. As shown in Table II, this helps the GS model converge faster and achieve higher reconstruction quality.
III-B Quantitative and qualitative evaluation.
We use standard metrics such as PSNR and SSIM for consistent and meaningful comparisons. Since our concern is the quality improvement of the object representation, evaluating the entire image would be less relevant. Therefore, scores are computed only on the pixels within the projected AABB.
| MethodsMetrics | PSNR | SSIM | #G | #G in box | Train |
| Baselines: 50K iters | |||||
| Full Scene images (1) | 21.23 | 0.917 | 3.72M | 8.73K | 98m |
| Full Object images (2) | 21.57 | 0.922 | 3.28M | 16.12K | 94m |
| Ours | |||||
| Scene-GS: 20K iters (3) | 20.78 | 0.911 | 3.02M | 6.99K | 35m |
| Object-GS: 30K iters (4) | 21.99 | 0.924 | 1.04M | 74.74K | 35m |
[…]
Table I presents quantitative comparisons between our method and baseline models. Our proposed composition approach, which combines Scene-GS (3) and Object-GS (4), achieves a higher ROI score than all baselines. As shown Figure 3, it produces noticeably sharper object details compared to baseline methods and Scene-GS alone. Table I also reports the total number of Gaussians (#G) and those within the ROI bounding box. In Object-GS (4), optimization is applied not only to Gaussians inside the box but also to outer ones to reduce floater artifacts. Notably, Object-GS (4) alone (boxed in red) achieves a higher PSNR than the composite model and all baselines. This demonstrates its strength in capturing fine details on the object. However, due to its limited scene coverage, overall image quality is reduced, as seen in rendering (b) of Figure 3. Row (c) shows renderings of another object, a decorative statue. Similar observations hold.
III-C Ablations.
| MethodsMetrics | PSNR | SSIM |
[…]
| No Scene-GS Init + 9-params GP (1) | 21.65 | 0.919 |
| No 3 static ROI params (6-params GP) (2) | 21.20 | 0.908 |
| No Advanced GP-based model (3) | 19.18 | 0.877 |
We evaluate the impact of Advanced Camera Selection and Scene-GS initialization for object training in Table II. In ablation setting (1): we use SfM point cloud initialization instead of trained Gaussians from Scene-GS, combined with Advanced 9-parameters GP camera selection. This results in a slight drop in ROI reconstruction quality, compared to the full method. Settings (2) and (3) both use Scene-GS initialization. (2) employs the original GP-based model without the three static ROI information, while (3) uses the simple selection based on static criteria only instead of the Advanced GP-based selection. The results highlight the effectiveness of the Advanced Selection approaches using GP model-based optimization. Moreover, incorporating additional object-specific information further improves the GP model.
IV Conclusion and Perspectives
We presented ROI-GS, a region-aware 3DGS method that enhances the level of detail for user-selected objects of interest by leveraging object-guided camera selection during training and a simple yet effective composition strategy during rendering. Our approach highlights the flexibility that it can enhance object of interest fidelity when applied to any global-optimization model, without compromising surrounding scene quality. While ROI-GS enables efficient and high-quality object view synthesis, several limitations remain.
First, our method relies on bounding boxes for object delineation, which can lead to imprecise boundaries. Incorporating object-level segmentation instead could improve spatial accuracy and reduce visual artifacts, particularly in cluttered or occluded areas.
Second, current LOD generation performed during GS training is static during rendering. Introducing dynamic LOD adjustment at render time based on the camera viewpoint could enable smarter resource allocation and better performance. Additionally, ROI-GS has a training time that scales linearly with the number of objects, and the size of the composed model depends on both the number and size of these objects. While a longer training time may be acceptable, an excessive number of Gaussians could negatively impact real-time rendering performance. In such cases, dynamic LOD becomes essential. For instance, for objects viewed from a distance, we could simply ignore the Object-GS composition and fall back to the Scene-GS representation.
Finally, our composition strategy may suffer from marginal inconsistencies near ROI boundaries. Gaussians centered inside an ROI may still extend into the background, and vice versa, causing slight overlaps between Object-GS and Scene-GS representations. While these artifacts are generally negligible in practice, better spatial filtering or blending strategies at the ROI boundary could further improve the compositional accuracy.